# Dynamical Edge Modes in $p$-form Gauge Theories

### v1

@@ 3.1 Phase space split
This vanishes whenever $\lambda$’s boundary pullback does, i.e. when $\lambda$ is a small gauge transformation, but if $i_{\partial\Sigma}\lambda\neq 0$ then this can be nonzero because the DEM boundary condition allows nonzero flux $i_{\partial\Sigma}*_{\Sigma}E$ through the boundary. We give a name to this flux, which is a $(p-1)$-form on $\partial\Sigma$, defining
@@
$*_{\partial\Sigma}E_{\perp}=i_{\partial\Sigma}*_{\Sigma}Sd\beta\;.$ (3.26)
@@ 3.2 Hamiltonian split
It can be understood as follows. Since $\beta$ is weakly $S$-harmonic, it is determined by its boundary data. Part of this data is $i_{\partial\Sigma}*_{\Sigma}\beta=0$, and for the other part we can take either “Dirichlet” data $i_{\partial\Sigma}\beta$ or “Neumann” data $E_{\perp}$ with $*_{\partial\Sigma}E_{\perp}=Si_{\partial\Sigma}*_{\Sigma}d\beta$ [63]. We then define the Dirichlet-to-Neumann operator $K:\Omega^{p-1}(\partial\Sigma)\to\Omega^{p-1}(\partial\Sigma)$ as the map from Dirichlet data to the corresponding Neumann data. This is once again a direct generalization of the Maxwell case. In terms of $K$ we can write
@@ 3.3 Horizon limit
The awkward inclusion of many upper metrics in the second equation is the cost of avoiding Christoffel symbols. We relegate most of the calculation to appendix B. There we expand in eigenmodes of the Laplacian $\Delta_{p-1}^{\partial\Sigma}$ for $(p-1)$-forms on $\partial\Sigma$ and we show that the mode coefficients obey a simple equation at leading order in small $r$. Explicitly, we write
@@ 4.1 Edge Partition Function
In this subsection we compute the thermal partition function in the special case where $\Sigma$ has the topology of a $(D-1)$-ball. Most of the topological factors then drop out. The most important example in this class of manifolds is the de Sitter static patch. We work in the horizon limit of section 3.3, so throughout this section we set
@@
For uniformity of discussion we also assume $22\leq p\leq D-$ 2, although most formulas apply to $p=0,1$ as well without modification. Upon quantization, the factorization of phase space implies a tensor factorization of Hilbert space,
@@
As discussed above, the bulk degrees of freedom are precisely those of the PMC boundary condition, so we can identify $Z_{\rm bulk}(\beta)=Z_{\rm PMC}(\beta)$. Our interest is in $Z_{\rm edge}(\beta)$. We will evaluate it as a canonical trace in the |$\ket{E_{\perp}}$ basis of edge states with definite $E_{\perp}$. There are no harmonic $(p-1)$-forms on $\partial\Sigma=S^{D-2}$ for $22\leq p\leq D-$ 2 so $K$ has no zero modes and $E_{\perp}$ can be any co-closed form on $\partial\Sigma$. Apparently also $i_{\partial\Sigma}\beta=K^{-1}E_{\perp}$ can be any co-closed form. The same argument implies that $i_{\partial\Sigma}\alpha$ can be any co-closed form. It will be convenient to expand in orthonormal eigenmodes of the Laplacian on $\partial\Sigma$,
@@
Here $\omega_{n}(x^{a})$ is a co-closed $(p-1)$-form on $\partial\Sigma$ satisfying
@@
We see that $\alpha_{n}$ and $E_{\perp,n}$ are conjugate variables, with each mode amounting to a copy of free 1D quantum mechanics. We define a state |$\ket{E_{\perp,n}}$ for each individual mode, normalized as
@@
as the size of the range of $\alpha_{n}$. The full state is
|$\ket{E_{\perp}}=\prod_{n}\ket{E_{\perp,n}}\;.$ (4.11)
@@
$\int\mathcal{D}E_{\perp}\ket{E_{\perp}}\bra{E_{\perp}}=\int\prod_{n}\frac{dE_{\perp,n}}{2\pi}\ket{E_{\perp,n}}\bra{E_{\perp,n}}.$ | . (4.12)
@@
$Z_{\rm edge}(\beta)$ $=\int\mathcal{D}E_{\perp}\bra{E_{\perp}}e^{-\beta H_{\rm edge}}\ket{E_{\perp}}$ (4.13)
@@
[…]
@@
$|\mathcal{G}^{\partial\Sigma}|$ $\equiv|\mathcal{G}^{\partial\Sigma}_{p-1}/(\mathcal{G}^{\partial\Sigma}_{p-2}/\dots)|=|\mathcal{G}^{\partial\Sigma}_{\rm nontriv}|\,|U(1)|^{(-)^{p+1}}\prod_{k=0}^{p-2}\left(\det^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k+1}}\;,$ (4.15)
@@
[…]
@@
$=\left(|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\left(\det^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k}}\right)\det\left(\frac{\log\varepsilon^{-1}}{\mu^{2}\kappa\beta}\Delta_{p-1}^{T,\partial\Sigma}\right)^{\frac{1}{2}}$
$=\left(\frac{\log\varepsilon^{-1}}{\kappa\beta}\right)^{\frac{1}{2}(-)^{p}+\rm{anomaly}}\det(\mu^{-2}\Delta_{p-1}^{T,\partial\Sigma})^{\frac{1}{2}}|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\left(\det^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k}}$
where the factor $\left(\frac{\log\varepsilon^{-1}}{\kappa\beta}\right)^{\frac{1}{2}(-)^{p}+\rm{anomaly}}$ comes from scaling the factor $\frac{\log\varepsilon^{-1}}{\kappa\beta}$ out of the $(p-1)$-form determinant. The general rule for pulling out a constant from a (full) $p$-form Laplacian $\Delta_{p}$ is [66]
$\det^{\prime}(a\Delta_{p})=a^{-\dim\ker\Delta_{p}+\rm{anomaly}}\det^{\prime}(\Delta_{p})\;,$ (4.18)
@@
$\det^{\prime}(a\Delta^{T}_{p})=a^{-\sum_{k=0}^{p}(-)^{p-k}b_{k}+\rm{anomaly}}\det^{\prime}(\Delta^{T}_{p})\;.$ (4.19)
In the present case where $\partial\Sigma$ is topologically a sphere, the only nonzero Betti numbers are $b_{0}=b_{D-2}=1$. Aside from this factor, we recognize the rest of the final expression in (4.17) as the reciprocal of the partition function $\mathcal{Z}_{\rm PI}^{p-1}[\partial\Sigma]$ of a $(p-1)$-form gauge theory on $\partial\Sigma$ with fundamental charge $q$, [^9] In general the fundamental charge of $p$-form gauge theory in $D$ dimensions has mass dimension $[q]=-\frac{1}{2}D+p+1$, so it is the same mass dimension as for $(p-1)$-form gauge theory in $D-2$ dimensions. so we can finally write
@@ 4.2.1 Comparing with $S^{D}$ with no hole
$\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]=$ $|U(1)|^{(-)^{p}}\left(\det^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{p}}{2}}\prod_{k=1}^{p}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (4.23)
@@
is the Harish-Chandra character for a massless $p$-form gauge field in $dS_{D}$. As pointed out in [2] and recalled in appendix C.1.1, (4.25) or more generally (C.9) can be independently defined in a $dS_{D}$ static patch at any inverse temperature $\beta$, with a Lorentzian interpretation as a “quasicanonical” trace over local propagating degrees of freedom in the static patch. The formula (4.25) corresponds to $\beta=2\pi$. In [14], the physical interpretation of the edge part $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was not as well understood, and $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was presented in terms of character integrals of the form (4.25) on a series of lower-dimensional de Sitter spaces, namely $dS_{D-2},dS_{D-4},\cdots,dS_{D-2p}$ (when $2p<D$). Later in [51], (the non-zero mode part of) $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was recognized to be the reciprocal of the partition function of a massless $(p-1)$-form gauge theory on $S^{D-2}$. In appendix C.4, we re-derive the split (4.24) following the approach in [2]. We show that including the dependence on the fundamental charge $q$ in (2.33), we have exactly
@@
We recall that $Z_{\rm bulk}(2\pi)=Z_{\rm PMC}(2\pi)$, i.e. it captures the bulk propagating degrees of freedom satisfying the PMC boundary condition; this is thus expected to agree with $\mathcal{Z}^{p}_{\rm bulk}\left[dS_{D}\right]$ in the $\varepsilon\to 0$ limit (up to possible non-universal terms). This expectation bears out in even dimensions, but in odd dimensions there is a slight discrepancy. [^12] For Maxwell in certain simple geometries $\log Z_{\rm PMC}$ can be recast as $D-2$ Neumann scalar partition functions, up to a term $\log\frac{Z_{D}}{Z_{N}}$, where $Z_{D(N)}$ is the scalar partition function with the Dirichlet (Neumann) boundary condition [10, 1]. This term is of order $\log\log\varepsilon^{-1}$, which in even $D$ will not affect the coefficient of $\log\mu R$. On the other hand, we have
@@ 4.3.1 Log-coefficients in even $D$
Using the prescription described above, one can compute the total log-coefficients of $\log\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]$ for arbitrary $20\leq p\leq D-$ 2 and even $D$. Here we display the explicit results up to $D=10$:
$D$ $p$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
$2$ $\frac{1}{3}$
$4$ $\frac{29}{90}$ $-\frac{31}{45}$ $\frac{209}{90}$
$6$ $\frac{1139}{3780}$ $-\frac{1271}{1890}$ $\frac{221}{210}$ $-\frac{5051}{1890}$ $\frac{16259}{3780}$
$8$ $\frac{32377}{113400}$ $-\frac{4021}{6300}$ $\frac{2603}{2520}$ $-\frac{8051}{5670}$ $\frac{7643}{2520}$ $-\frac{29221}{6300}$ $\frac{712777}{113400}$
$10$ $\frac{2046263}{7484400}$ $-\frac{456569}{748440}$ $\frac{13228}{13365}$ $-\frac{5233531}{3742200}$ $\frac{1339661}{748440}$ $-\frac{12717931}{3742200}$ $\frac{66688}{13365}$ $-\frac{4947209}{748440}$ $\frac{61921463}{7484400}$
@@
$D$ $p$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
$4$ $\frac{29}{90}$ $-\frac{16}{45}$ $\frac{29}{90}$
$6$ $\frac{1139}{3780}$ $-\frac{331}{945}$ $\frac{229}{630}$ $-\frac{331}{945}$ $\frac{1139}{3780}$
$8$ $\frac{32377}{113400}$ $-\frac{1592}{4725}$ $\frac{545}{1512}$ $-\frac{1042}{2835}$ $\frac{545}{1512}$ $-\frac{1592}{4725}$ $\frac{32377}{113400}$
$10$ $\frac{2046263}{7484400}$ $-\frac{303601}{935550}$ $\frac{657683}{1871100}$ $-\frac{342019}{935550}$ $\frac{276929}{748440}$ $-\frac{342019}{935550}$ $\frac{657683}{1871100}$ $-\frac{303601}{935550}$ $\frac{2046263}{7484400}$

### v2

@@ 3.1 Phase space split
This vanishes whenever $\lambda$’s boundary pullback does, i.e. when $\lambda$ is a small gauge transformation, but if $i_{\partial\Sigma}\lambda\neq 0$ then this can be nonzero because the DEM boundary condition allows nonzero flux $i_{\partial\Sigma}*_{\Sigma}E$ through the boundary. We give a name to this flux, which is a (p1)(p-1)-form on $\partial\Sigma$, defining
@@
$_{\partial\Sigma}E_{\perp}=i_{\partial\Sigma}*_{\Sigma}Sd\beta\;.$ (3.26)
@@ 3.2 Hamiltonian split
It can be understood as follows. Since $\beta$ is weakly $S$-harmonic, it is determined by its boundary data. Part of this data is $i_{\partial\Sigma}*_{\Sigma}\beta=0$, and for the other part we can take either “Dirichlet” data $i_{\partial\Sigma}\beta$ or “Neumann” data $E_{\perp}$ with $_{\partial\Sigma}E_{\perp}=Si_{\partial\Sigma}*_{\Sigma}d\beta$ [66]. We then define the Dirichlet-to-Neumann operator $K:\Omega^{p-1}(\partial\Sigma)\to\Omega^{p-1}(\partial\Sigma)$ as the map from Dirichlet data to the corresponding Neumann data. This is once again a direct generalization of the Maxwell case. In terms of $K$ we can write
@@ 3.3 Horizon limit
The awkward inclusion of many upper metrics in the second equation is the cost of avoiding Christoffel symbols. We relegate most of the calculation to appendix B. There we expand in eigenmodes of the Laplacian $\Delta_{p-1}^{\partial\Sigma}$ for (p1)(p-1)-forms on $\partial\Sigma$ and we show that the mode coefficients obey a simple equation at leading order in small $r$. Explicitly, we write
@@ 4.1 Edge Partition Function
In this subsection we compute the thermal partition function in the special case where $\Sigma$ has the topology of a (D1)(D-1)-ball. Most of the topological factors then drop out. The most important example in this class of manifolds is the de Sitter static patch. We work in the horizon limit of section 3.3, so throughout this section we set
@@
For uniformity of discussion we also assume $2\leq p\leq D-2$, although most formulas apply to $p=0,1$ as well without modification. Upon quantization, the factorization of phase space implies a tensor factorization of Hilbert space,
@@
[…]
@@
Here $\omega_{n}(x^{a})$ is a co-closed (p1)(p-1)-form on $\partial\Sigma$ satisfying
@@
[…]
@@
$\langle\,E_{\perp,n}\mid E_{\perp,n}\,\rangle=2\pi\delta(0)=\int_{-\infty}^{\infty}d\alpha_{n}$ (4.10)
as the size of the range of $\alpha_{n}$. The full state is
@@
[…]
@@
[…]
@@
$=\left(\prod_{n}\frac{\mu}{\sqrt{2\pi}}\int_{-\infty}^{\infty}d\alpha_{n}\right)\left(\prod_{n}\frac{\lambda_{n}\log\varepsilon^{-1}}{\mu^{2}\kappa\beta}\right)^{\frac{1}{2}}$
@@
$|\mathcal{G}^{\partial\Sigma}|\equiv|\mathcal{G}^{\partial\Sigma}_{p-1}/(\mathcal{G}^{\partial\Sigma}_{p-2}/\dots)|=|\mathcal{G}^{\partial\Sigma}_{\rm nontriv}|\,|U(1)|^{(-)^{p+1}}\prod_{k=0}^{p-2}\left(\det\hskip 0.0pt^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k+1}}\;,$ (4.15)
@@
$|U(1)|=\int_{0}^{\frac{2\pi}{q}\sqrt{V_{\partial\Sigma}}}\frac{\mu^{p}}{\sqrt{2\pi}}d\phi_{0}=\frac{\mu^{p}}{q}\sqrt{2\pi V_{\partial\Sigma}}\;.$ (4.16)
@@
$=\left(|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\left(\det\hskip 0.0pt^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k}}\right)\det\left(\frac{\log\varepsilon^{-1}}{\mu^{2}\kappa\beta}\Delta_{p-1}^{T,\partial\Sigma}\right)^{\frac{1}{2}}$
$=\left(\frac{\log\varepsilon^{-1}}{\kappa\beta}\right)^{\frac{1}{2}(-)^{p}+\rm{anomaly}}\det(\mu^{-2}\Delta_{p-1}^{T,\partial\Sigma})^{\frac{1}{2}}|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\left(\det\hskip 0.0pt^{\prime}(\mu^{-2}\Delta_{k}^{T,\partial\Sigma})^{-\frac{1}{2}}\right)^{(-)^{p-k}}$
where the factor $\left(\frac{\log\varepsilon^{-1}}{\kappa\beta}\right)^{\frac{1}{2}(-)^{p}+\rm{anomaly}}$ comes from scaling the factor $\frac{\log\varepsilon^{-1}}{\kappa\beta}$ out of the (p1)(p-1)-form determinant. The general rule for pulling out a constant from a (full) $p$-form Laplacian $\Delta_{p}$ is [69]
$\det\hskip 0.0pt^{\prime}(a\Delta_{p})=a^{-\dim\ker\Delta_{p}+\rm{anomaly}}\det\hskip 0.0pt^{\prime}(\Delta_{p})\;,$ (4.18)
@@
$\det\hskip 0.0pt^{\prime}(a\Delta^{T}_{p})=a^{-\sum_{k=0}^{p}(-)^{p-k}b_{k}+\rm{anomaly}}\det\hskip 0.0pt^{\prime}(\Delta^{T}_{p})\;.$ (4.19)
In the present case where $\partial\Sigma$ is topologically a sphere, the only nonzero Betti numbers are $b_{0}=b_{D-2}=1$. Aside from this factor, we recognize the rest of the final expression in (4.17) as the reciprocal of the partition function $\mathcal{Z}_{\rm PI}^{p-1}[\partial\Sigma]$ of a (p1)(p-1)-form gauge theory on $\partial\Sigma$ with fundamental charge qq, [^9] In general the fundamental charge of $p$-form gauge theory in $D$ dimensions has mass dimension $q]=-\frac{1}{2}D+p+1$, so it is the same mass dimension as for (p1)(p-1)-form gauge theory in D2D-2 dimensions. so we can finally write
@@ 4.2.1 Comparing with $S^{D}$ with no hole
$\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]=$ $|U(1)|^{(-)^{p}}\left(\det\hskip 0.0pt^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{p}}{2}}\prod_{k=1}^{p}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (4.23)
@@
is the Harish-Chandra character for a massless $p$-form gauge field in $dS_{D}$. As pointed out in [2] and recalled in appendix C.1.1, (4.25) or more generally (C.9) can be independently defined in a $dS_{D}$ static patch at any inverse temperature $\beta$, with a Lorentzian interpretation as a “quasicanonical” trace over local propagating degrees of freedom in the static patch. The formula (4.25) corresponds to $\beta=2\pi$. In [14], the physical interpretation of the edge part $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was not as well understood, and $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was presented in terms of character integrals of the form (4.25) on a series of lower-dimensional de Sitter spaces, namely $dS_{D-2},dS_{D-4},\cdots,dS_{D-2p}$ (when $2p<D$). Later in [53], (the non-zero mode part of) $\mathcal{Z}^{p}_{\rm edge}\left[S^{D}\right]$ was recognized to be the reciprocal of the partition function of a massless (p1)(p-1)-form gauge theory on $S^{D-2}$. In appendix C.4, we re-derive the split (4.24) following the approach in [2]. We show that including the dependence on the fundamental charge $q$ in (2.33), we have exactly
@@
We recall that $Z_{\rm bulk}(2\pi)=Z_{\rm PMC}(2\pi)$, i.e. it captures the bulk propagating degrees of freedom satisfying the PMC boundary condition; this is thus expected to agree with $\mathcal{Z}^{p}_{\rm bulk}\left[dS_{D}\right]$ in the $\varepsilon\to 0$ limit (up to possible non-universal terms). This expectation bears out in even dimensions, but in odd dimensions there is a slight discrepancy. [^12] For Maxwell in certain simple geometries $\log Z_{\rm PMC}$ can be recast as D2D-2 Neumann scalar partition functions, up to a term $\log\frac{Z_{D}}{Z_{N}}$, where $Z_{D(N)}$ is the scalar partition function with the Dirichlet (Neumann) boundary condition [10, 1]. This term is of order $\log\log\varepsilon^{-1}$, which in even $D$ will not affect the coefficient of $\log\mu R$. On the other hand, we have
@@
In addition to its technical tractability, this $S^{D}$ example is important because of its relevance to de Sitter space quantum gravity. According to the Gibbons-Hawking proposal [75], $S^{D}$ partition functions compute quantum corrections to the de Sitter horizon entropy, which provide precision tests for candidate microscopic models in the spirit of [76, 77, 78, 79, 80, 81, 82, 83, 84, 2]. As shown in [2], any spinning 1-loop $S^{D}$ partition functions admit a bulk-edge split as in (4.24). Clarifying the physical meaning of each part may help identify specific sectors of candidate microscopic models. The discussion above establishes that the edge partition function can be interpreted as a thermal partition function of the edge mode sector, corresponding to large gauge transformations residing on the stretched horizon. This provides a concrete physical picture of the edge sector, which was not fully understood in previous studies.
@@ 4.3.1 Log-coefficients in even $D$
Using the prescription described above, one can compute the total log-coefficients of $\log\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]$ for arbitrary $0\leq p\leq D-2$ and even $D$. Here we display the explicit results up to $D=10$:
$D$ pp 0 11 22 33 44 55 66 77 88
22 $\frac{1}{3}$
44 $\frac{29}{90}$ $\frac{31}{45}$ $\frac{209}{90}$
66 $\frac{1139}{3780}$ $\frac{1271}{1890}$ $\frac{221}{210}$ $\frac{5051}{1890}$ $\frac{16259}{3780}$
88 $\frac{32377}{113400}$ $\frac{4021}{6300}$ $\frac{2603}{2520}$ $\frac{8051}{5670}$ $\frac{7643}{2520}$ $\frac{29221}{6300}$ $\frac{712777}{113400}$
1010 $\frac{2046263}{7484400}$ $\frac{456569}{748440}$ $\frac{13228}{13365}$ $\frac{5233531}{3742200}$ $\frac{1339661}{748440}$ $\frac{12717931}{3742200}$ $\frac{66688}{13365}$ $\frac{4947209}{748440}$ $\frac{61921463}{7484400}$
@@ 66 $\frac{1139}{3780}$ $\frac{1271}{1890}$ $\frac{221}{210}$ $\frac{5051}{1890}$ $\frac{16259}{3780}$
$D$ pp 0 11 22 33 44 55 66 77 88
44 $\frac{29}{90}$ $\frac{16}{45}$ $\frac{29}{90}$
66 $\frac{1139}{3780}$ $\frac{331}{945}$ $\frac{229}{630}$ $\frac{331}{945}$ $\frac{1139}{3780}$
88 $\frac{32377}{113400}$ $\frac{1592}{4725}$ $\frac{545}{1512}$ $\frac{1042}{2835}$ $\frac{545}{1512}$ $\frac{1592}{4725}$ $\frac{32377}{113400}$
1010 $\frac{2046263}{7484400}$ $\frac{303601}{935550}$ $\frac{657683}{1871100}$ $\frac{342019}{935550}$ $\frac{276929}{748440}$ $\frac{342019}{935550}$ $\frac{657683}{1871100}$ $\frac{303601}{935550}$ $\frac{2046263}{7484400}$